# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v1

@@ 8 CONCLUSION AND FUTURE WORK
We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes. This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics. Additionally, we incorporated a compact Poisson filter into the fluid dynamics solver to achieve efficient, high-performance simulations with a balance of stability and computational accuracy. The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes. Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both cloud-to-ground and intra-cloud lightning. Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
@@
Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena. Enhancements to thunderstorm microphysics, such as more detailed electrification processes and turbulence modeling, will improve the realism of lightning generation. The framework will also be extended to simulate additional thunderstorm types, including Mesoscale Convective Complex (MCC), Mesoscale Convective Vortex (MCV), and derechos, broadening its coverage of mesoscale systems. Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models. Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and interactive simulations.

### v3

@@ 8. CONCLUSION AND FUTURE WORK
We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes. This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics. The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes. Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both CG and IC lightning. Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
Figure 9. Comparison of cloud coverage evolution (top row) ,cloud fraction (middle row) and lightning activity evolution(bottom row)across four distinct regions: Florida, New Mexico, Japan, and California. The figures showcase the distinct patterns of atmospheric dynamics influenced by regional environmental conditions.
@@
Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena. Enhancements to thunderstorm microphysics, such as more detailed electrification processes and turbulence modeling, will improve the realism of lightning generation. The framework will also be extended to simulate additional thunderstorm types, including Mesoscale Convective Complex (MCC), Mesoscale Convective Vortex (MCV), and derechos, broadening its coverage of mesoscale systems. Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models. Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and efficient simulations.